Skip to Search

 
#

Support Services Meeting Notes
March 16, 2009

Present: Bill Covey, Will Chaney, Lori Driscoll, Rob Roberts, Laurie Taylor
Facilities
  1. Dehumidifier in the IT Server room is still out.
  2. Incident report last night: man found with wad of toilet paper to his head. EMTs were called. Police took a head print off the door.
  3. 4th floor East water damage - not recent - sent to PPD.
  4. Set points for DLC: 72-77 degrees. Temperature in the room to be kept within the set points. Rob spoke with john Lawson about the room.
  5. Maps is done. The stairwell is back together.
  6. Education - alarms - Iona went out a door that set off an alarm and then called PPD to cancel it. They hadn't received notification. The reason why is because they only get the notification after midnight.
  7. HSC - Vern sent Rob an email for three more doors to be covered.
  8. Suncom accounts - Rob turned it over to Grace in Fiscal Services. Setting up conference calls put them all in Lori Driscoll's name. Now all accounts will appear to Fiscal Services by Departmental name.
Access Services
  1. Lori met with Jan Swanbeck and some other folks - we are arranging for Documents and Maps to have regular circulation periods like the rest of the departments. They are also going to be in the campus delivery program with the Law School.
  2. Delivery program is going well. They are starting to let folks know over in Law.
  3. New Oak Hammock form.
  4. Launched the new Borrower page. It has different information now. There were complaints from branches that it was hard to get to from the main page.
  5. Kelli started working with us. She is doing great. We will still hire another half-time person for this position.
  6. Working on evaluations.
IT
  1. Amy has been working on the Green Team's request to shut down monitors after 10 minutes of inactivity. She has the policy working. Bill suggests trying his machine to test it.
  2. EPO server - issue with monitoring. Will is going to spend time with Bobbie on this issue.
  3. DLC equipment request - waiting for Dell quote; got one from GovConnection.
  4. DLC - Logan meeting with them for solutions.
DLC
  1. Space is our top concern at the moment. Right now, we are just getting by on work-arounds. Space monitoring is not working. Logan is working on solutions. The reporting tool should be available for everyone. Need JPG2K3 space. Laurie to work on permission levels.
  2. FCLA - new service on the catalog that shows digital holdings - it's only showing PALMM mark files. Have to redo it for the rest of the collection.
  3. FAMU - asked us to pull our archival files because FDA can't pull them fast enough.
  4. HSC - archivist on Friday - drives to the Cade papers - 900 newspapers - we are getting the drives.
  5. Oak Hammock presentation went well. They liked the forms. Auto-renew? Expiration date is set way in the future. They are very happy with the Library services and liked the DVD checkout. They have their own library - no, it's not possible to merge it with ours. One of the residents, Dr. Fatina, approached us in digitizing letters from the Dresden opera.
  6. Grants for Aerials and three Oral Histories were submitted.
Directors
  1. Directory front - affiliation for the Library - we've asked for roles to control what patrons have the affiliation. The affiliation would be handed out on the desk. Should it be available in all branches, or just in West? Winston is looking into how to get it working. They are still writing the roles to manage it. There is light at the end of the tunnel.
  2. Next town meeting on the budget - no new info. Anti-climatic.
  3. Bill and Michelle to return to their positions, as announced by Judy. Rachel is scheduling a meeting to get together with all of her head folks on April 9th. Leave room in your schedules.
  4. HSC integration - photocopy contract may be the issues that pushes the separate IT departments together.

 

University of Florida Home Page